Husain, Inc.’s income statement and other financial information for the current year is presented below.

Hussain, Inc.
Income Statement
For the year ended December 31
Sales revenue $ 199,131
Cost of goods sold 72,360
Gross profit 126,771
Selling, general and administrative expenses 13,385
Operating income 113,386
Interest expense 4,047
Income before taxes 109,339
Income tax expense 4,214
Net income $ 105,125

Balance sheet information:
Current assets $ 254,000
Noncurrent assets 508,000
Current liabilities 58,000
Long-term debt 116,000


Required:

  1. Perform vertical analysis of the income statement.
  1. Calculate the debt-to-assets ratio.
  1. Calculate the times interest earned ratio.nss. nc.s income stetemen and other triancel itomstion lor the current year presented beow. Hussain, Ic Income Ststenent For
Answer #1

To Condust Vertical analysis of INCOME STATEMENT Sales figure is Considered as base figure and all other components of income statement like cost of sales , operating expenses gross profit income taxes and net income etc are shown as percentage of sales

In vertical analysis the percentage is computed by using the following formula

Percentage Base =(amount of individual item/Amount of base)X100

Therefore Sales base figure is $199131

HUSSAIN INC

INCOME STATEMENT FOR THE YEAR ENDED DECEMBER 31
PARTICULARS AMOUNT $ PERCENTAGE
Sales 199 131
Cost of goods Sold 72360 =(72360/199131)X100=36.33%
Gross Profit 126 771 =(126 771/ 199 131)X100 =63.66%
Selling General &Admin Expenses 13385 =(13385/199131)X100=6.72%
Oparating Income 113386 =(113386/199131)X100=56.94%
Interest Expenses 4047 =(4047/199131)X100=2.03%
Incoem Befroe Taxes 109339 =(109339/199131)X100=54.90%
Income Tax Expenses 4214 =(4214/199131)X100=2.11%
Net Income 105125 =(105125/199131)X10=52.79%

Debt to Assets Ratio=Short Term Debt +Long term Debt/Total Assets

Long term Debt $116000

Total Assets $762000

Debt to Assets Ratio =$116000/$762000=0.15 Times

Times Interest Earned ratio is calculated by dividing net income before interest and income tax by interest expenses

Times Interest Earned ratio= Income before interest and income taxes or EBIT/Interest Expenses

Times Interest Earned ratio =$113386/$4047=28.01 Times

Thanks